OpenStudy (anonymous):

a 10 foot piece of string is cut into two pieces so that the longer piece is 1 foot longer than twice the shorter piece. If the shorter piece is X feet long. Find the lengths of both pieces

OpenStudy (albert0898):

Short Piece = x Long Piece = 2x + 1 (2x + 1) + x = 10 3x + 1 = 10 3x = 9 x = 3 Short Piece = x = 3 Long Piece = 2x + 1 = 7 Do you understand or do you need me to explain?

OpenStudy (anonymous):

gotcha! thank you